# Scalable Valuation of Human Feedback through Provably Robust Model Alignment

### v1

@@ E.2 Proof for Corollary 2
Suppose that the policy gradient $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ is bounded by $C$ and satisfies $L$-Lipchitz in $\theta$, where $0<C<\infty$ and $0<L<\infty$. Then, under Theorem 2, the IF of DPO does not satisfy the redescending property in Definition 2, i.e., $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{DPO}}(x,\theta,p_{\mathcal{D}})\neq 0$.
@@
From the positive definite assumption in Theorem 2 and the $L$-Lipschitsness assumption of $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ in $\theta$, we can see that $F_{\theta^{*}}(s)$ is $(\sigma(-g_{\theta^{*}}(s))\cdot L)$-Lipschits continuous in $\theta$. We thus have $0<\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}(s)]\leq L<\infty$, where we use Lemma 1.2.2 in Nesterov [75] and the fact that $\sigma(-g_{\theta}(s))\leq 1$. Furthermore, from the assumption that $\|\nabla_{\theta}\log\pi_{\theta}(y\mid x)\|\leq C$ ($0<C<\infty$), we have $\|\nabla_{\theta}\log\pi_{\theta}(y_{\text{win}}\mid x)-\nabla_{\theta}\log\pi_{\theta}(y_{\text{lose}}\mid x)\|\leq 2C$ from the triangle inequality. Thus, the term $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{DPO}}(x,\theta,p_{\mathcal{D}})$ can be upper bounded as $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{DPO}}(x,\theta,p_{\mathcal{D}})\leq 2C/L^{\prime}$, where $0<L^{\prime}\leq\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}(s)]$ and we use the fact that $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathbb{E}_{p(s_{\textrm{flip}})}[\sigma(-g_{\theta^{*}}(s_{\textrm{flip}}))]=\mathbb{E}_{p(s_{\textrm{flip}})}[\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\sigma(-g_{\theta^{*}}(s_{\textrm{flip}}))]=1$, where we use the bounded convergence theorem. The fact $0<2C/L^{\prime}<\infty$ according to $0<C<\infty$ and $0<L^{\prime}<\infty$ completes the proof. ∎
@@ E.3 Proof for Theorem 4
We next evaluate $\nabla_{\theta}A(\theta)=\nabla_{\theta}\mathbb{E}_{p_{\mathcal{D}}}[\sigma(g_{\theta}(s))^{\gamma}]$. Since $p_{\mathcal{D}}$ does not depend on $\theta$, from the cain rule, we can see $\nabla_{\theta}\mathbb{E}_{p_{\mathcal{D}}}[\sigma(g_{\theta}(s))^{\gamma}]=\mathbb{E}_{p_{\mathcal{D}}}[\nabla_{\theta}\sigma(g_{\theta}(s))^{\gamma}]$, where
@@
Furthermore, we show the following lemmas to precisely derive the IF for Hölder-DPO.
@@
Suppose that $\theta^{*}=\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}(p_{\mathcal{D}}\|\sigma(g_{\theta}))$. Let $0<\sigma(g_{\theta}(s))\leq 1$ for all $s\in\mathcal{D}$ and $0<\gamma<\infty$. Then, $\phi(h_{\theta^{*}})\neq 0$ and $\phi^{\prime}(h_{\theta^{*}})\neq 0$.
@@
From the definition of $\phi$ in Definition 3, we have $\phi(h_{\theta^{*}})\geq-h_{\theta^{*}}^{1+\gamma}$. Therefore, if $h_{\theta^{*}}\neq 0$, we can see that $\phi(h_{\theta^{*}})\neq 0$. Recall the definition of $h_{\theta^{*}}$ in Proposition 2, this condition holds when $\mathbb{E}_{p_{\mathcal{D}}}[\sigma(g_{\theta^{*}}(s))^{\gamma}]\neq 0$, which is satisfied under the following mild condition $0<\sigma(g_{\theta}(s))$ and $0<\gamma<\infty$. We further obtain $\phi^{\prime}(h_{\theta^{*}})\geq-(1+\gamma)h_{\theta^{*}}^{\gamma}$. This also shows that $\phi^{\prime}(h_{\theta^{*}})\neq 0$ when $0<\sigma(g_{\theta}(s))$ and $0<\gamma<\infty$ are satisfied. ∎
@@
Suppose that $\theta^{*}=\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}(p_{\mathcal{D}}\|\sigma(g_{\theta}))$. Let $0<\gamma<\infty$, and let $0\leq\sigma(g_{\theta}(s))$. Then, the first-order optimal condition of Hölder-DPO, $0=\nabla_{\theta}S_{\gamma}(p_{\mathcal{D}}\|\sigma(g_{\theta}))|_{\theta=\theta^{*}}$, holds if and only if $\mathbb{E}_{p_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}}(s)]=0$ and $\int F^{(1+\gamma)}_{\theta^{*}}(s)\mathrm{d}s=0$, where $F^{(\gamma)}_{\theta^{*}}(s)\coloneqq\sigma(g_{\theta}(s))^{\gamma}(1-\sigma(g_{\theta}(s)))\nabla_{\theta}g_{\theta}(s)$.
@@
The statement obviously holds when $\sigma(g_{\theta}(s))=0$; therefore, we hereafter consider the case when $0<\sigma(g_{\theta}(s))$. From the definition, we have
00 0 $=\nabla_{\theta}S_{\gamma}(p_{\mathcal{D}}\|\sigma(g_{\theta}))\bigg{|}_{\theta=\theta^{*}}$
$=\gamma\phi^{\prime}(h_{\theta^{*}})\mathbb{E}_{p_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}}(s)]+(1+\gamma)\left(\phi(h_{\theta^{*}})-\phi^{\prime}(h_{\theta^{*}})h_{\theta^{*}}\right)\bigg{(}\int F^{(1+\gamma)}_{\theta^{*}}(s)\mathrm{d}s\bigg{)}.$
This condition obviously holds when $\phi^{\prime}(h_{\theta^{*}})\mathbb{E}_{p_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}}(s)]=0$ and $\left(\phi(h_{\theta^{*}})-\phi^{\prime}(h_{\theta^{*}})h_{\theta^{*}}\right)\bigg{(}\int F^{(1+\gamma)}_{\theta^{*}}(s)\mathrm{d}s\bigg{)}=0$. From Lemma 1, we can see that $\phi(h_{\theta^{*}})\neq 0$ and $\phi^{\prime}(h_{\theta^{*}})\neq 0$. Therefore, this condition can be rewritten as $\mathbb{E}_{p_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}}(s)]=0$ and $\int F^{(1+\gamma)}_{\theta^{*}}(s)\mathrm{d}s=0$. This is also a necessary condition, which becomes immediately clear when we consider that $\nabla_{\theta}S_{\gamma}(p_{\mathcal{D}}\|\sigma(g_{\theta}))\bigg{|}_{\theta=\theta^{*}}$ is expressed as a linear combination and that conditions, $\phi(h_{\theta^{*}})\neq 0$ and $\phi^{\prime}(h_{\theta^{*}})\neq 0$, hold. This completes the proof. ∎
@@
$\nabla_{\theta}S_{\gamma}(\tilde{p}^{(\epsilon)}_{\mathcal{D}}\|\sigma(g_{\theta}))=\gamma\phi^{\prime}(\tilde{h}_{\theta})\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[F^{(\gamma)}_{\theta}(\tilde{s})]+(1+\gamma)\cdot\bigg{(}\phi(\tilde{h}_{\theta})-\tilde{h}_{\theta}\cdot\phi^{\prime}(\tilde{h}_{\theta})\bigg{)}\cdot\bigg{(}\int F^{(1+\gamma)}_{\theta}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)},$
where $F^{(\gamma)}_{\theta}(\tilde{s})\coloneqq\sigma^{\gamma}(g_{\theta}(\tilde{s}))\left(1-\sigma(g_{\theta}(\tilde{s}))\right)\nabla_{\theta}g_{\theta}(\tilde{s})$, $F^{(1+\gamma)}_{\theta}(\tilde{s})\coloneqq\sigma(g_{\theta}(\tilde{s}))\cdot F^{(\gamma)}_{\theta}(\tilde{s})$, $\tilde{h}_{\theta}\coloneqq\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[\sigma(g_{\theta}(\tilde{s}))^{\gamma}]/(\int\sigma(g_{\theta}(\tilde{s}))^{1+\gamma}\mathrm{d}\tilde{s})$, and $\tilde{s}=\{\tilde{x},\tilde{y}_{\textrm{win}},\tilde{y}_{\textrm{lose}}\}\sim\tilde{p}_{\mathcal{D}}^{(\epsilon)}$.
@@
00 0 $=\nabla_{\theta}S_{\gamma}(\tilde{p}^{(\epsilon)}_{\mathcal{D}}\|\sigma(g_{\theta}))\bigg{|}_{\theta=\theta^{*}(\epsilon)}$
$=\gamma\phi^{\prime}(\tilde{h}_{\theta^{*}(\epsilon)})\mathbb{E}_{\tilde{p}^{(\epsilon)}_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})]+(1+\gamma)\left(\phi(\tilde{h}_{\theta^{*}(\epsilon)})-\phi^{\prime}(\tilde{h}_{\theta^{*}(\epsilon)})\tilde{h}_{\theta^{*}(\epsilon)}\right)\bigg{(}\int F^{(1+\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)},$
where $F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})\coloneqq\sigma(g_{\theta^{*}(\epsilon)}(\tilde{s}))^{\gamma}\left(1-\sigma(g_{\theta^{*}(\epsilon)}(\tilde{s}))\right)\{\nabla_{\theta}g_{\theta}(\tilde{s})|_{\theta=\theta^{*}(\epsilon)}\}$ and $F^{(1+\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})\coloneqq\sigma(g_{\theta^{*}(\epsilon)}(\tilde{s}))\cdot F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})$.
@@
00 0 $=\gamma\bigg{\{}\underbrace{\frac{\partial}{\partial\epsilon}\phi^{\prime}(\tilde{h}_{\theta^{*}(\epsilon)})\mathbb{E}_{\tilde{p}^{(\epsilon)}_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})]}_{\mathrm{(I)}}\bigg{\}}$
@@
$\frac{\partial}{\partial\epsilon}\phi^{\prime}(h_{\theta^{*}(\epsilon)})\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})]$
$=\phi^{\prime\prime}(h_{\theta^{*}(\epsilon)})\cdot\frac{\partial h_{\theta^{*}(\epsilon)}}{\partial\epsilon}\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})]$
$\quad\quad+\phi^{\prime}(h_{\theta^{*}(\epsilon)})\bigg{\{}\int\bigg{\{}\frac{\partial}{\partial\epsilon}p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})\bigg{\}}F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})\mathrm{d}\tilde{s}+\mathbb{E}_{\tilde{p}_{\mathcal{D}}^{(\epsilon)}}\bigg{[}\frac{\partial}{\partial\epsilon}F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})\bigg{]}\bigg{\}}$
$=\phi^{\prime\prime}(h_{\theta^{*}(\epsilon)})\cdot\frac{\partial h_{\theta^{*}(\epsilon)}}{\partial\epsilon}\cdot\mathbb{E}_{\tilde{p}^{(\epsilon)}_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})]$
@@
$=\phi^{\prime\prime}(h_{\theta^{*}(\epsilon)})\cdot\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}\cdot\frac{\partial h_{\theta^{*}(\epsilon)}}{\partial\theta^{*}(\epsilon)}\cdot\mathbb{E}_{\tilde{p}^{(\epsilon)}_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})]$
$\quad\quad+\phi^{\prime}(h_{\theta^{*}(\epsilon)})\bigg{\{}\mathbb{E}_{p(s_{\textrm{flip}})}[F^{(\gamma)}_{\theta^{*}(\epsilon)}(s_{\textrm{flip}})]-\mathbb{E}_{p_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}(\epsilon)}(s)]+\mathbb{E}_{\tilde{p}_{\mathcal{D}}^{(\epsilon)}}\bigg{[}\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}\cdot H_{\theta^{*}(\epsilon)}^{(\gamma)}(\tilde{s})\bigg{]}\bigg{\}}$
$=\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}\bigg{\{}\phi^{\prime\prime}(h_{\theta^{*}(\epsilon)})\cdot\frac{\partial h_{\theta^{*}(\epsilon)}}{\partial\theta^{*}(\epsilon)}\cdot\mathbb{E}_{\tilde{p}^{(\epsilon)}_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})]+\phi^{\prime}(h_{\theta^{*}(\epsilon)})\cdot\mathbb{E}_{\tilde{p}_{\mathcal{D}}^{(\epsilon)}}\bigg{[}H_{\theta^{*}(\epsilon)}^{(\gamma)}(\tilde{s})\bigg{]}\bigg{\}}$
@@
$\frac{\partial}{\partial\epsilon}$ $\left(\phi(h_{\theta^{*}(\epsilon)})-\phi^{\prime}(h_{\theta^{*}(\epsilon)})h_{\theta^{*}(\epsilon)}\right)\bigg{(}\int F^{(1+\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)}$
=$=\bigg{\{}\frac{\partial}{\partial\epsilon}\left(\phi(h_{\theta^{*}(\epsilon)})-\phi^{\prime}(h_{\theta^{*}(\epsilon)})h_{\theta^{*}(\epsilon)}\right)\bigg{\}}\bigg{(}\int F^{(1+\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)}$
$\quad\quad\quad\quad\quad\quad\quad\quad+\left(\phi(h_{\theta^{*}(\epsilon)})-\phi^{\prime}(h_{\theta^{*}(\epsilon)})h_{\theta^{*}(\epsilon)}\right)\bigg{(}\int\frac{\partial}{\partial\epsilon}F^{(1+\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)}$
$=\left(-\phi^{\prime\prime}(h_{\theta^{*}(\epsilon)})\cdot\frac{\partial h_{\theta^{*}(\epsilon)}}{\partial\epsilon}\cdot h_{\theta^{*}(\epsilon)}\right)\bigg{(}\int F^{(1+\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)}$
$\quad\quad\quad\quad\quad\quad\quad\quad+\left(\phi(h_{\theta^{*}(\epsilon)})-\phi^{\prime}(h_{\theta^{*}(\epsilon)})h_{\theta^{*}(\epsilon)}\right)\bigg{(}\int\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}H_{\theta^{*}(\epsilon)}^{(1+\gamma)}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)}$
$=\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}\bigg{\{}\left(-\phi^{\prime\prime}(h_{\theta^{*}(\epsilon)})\cdot\frac{\partial h_{\theta^{*}(\epsilon)}}{\partial\theta^{*}(\epsilon)}\cdot h_{\theta^{*}(\epsilon)}\right)\bigg{(}\int F^{(1+\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)}$
$\quad\quad\quad\quad\quad\quad\quad\quad+\left(\phi(h_{\theta^{*}(\epsilon)})-\phi^{\prime}(h_{\theta^{*}(\epsilon)})h_{\theta^{*}(\epsilon)}\right)\bigg{(}\int H_{\theta^{*}(\epsilon)}^{(1+\gamma)}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)}\bigg{\}},$
@@
This completes the proof. ∎

### v2

@@ E.2 Proof for Theorem 3 (The Case of DPO)
Suppose that the policy gradient $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ is bounded by $C$ and satisfies $L$-Lipchitz in $\theta$, where $0<C<\infty$ and $0<L<\infty$. Then, under Theorem 2, the IF of DPO does not satisfy the redescending property in Definition 2, i.e., $\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{DPO}}(x,\theta,p_{\mathcal{D}})\|\neq 0$.
@@
From the positive definite assumption on the Hessian in Theorem 5 and the $L$-Lipschitz assumption on the gradient, it follows that $\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}(s)]$ is a positive definite matrix. Let $L^{\prime}=\lambda_{\mathrm{min}}(\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}(s)])>0$ be its minimum eigenvalue. Then the norm of its inverse is bounded: $\|(\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}(s)])^{-1}\|=1/L^{\prime}$. Furthermore, from the assumption that $\|\nabla_{\theta}\log\pi_{\theta}(y\mid x)\|\leq C$ ($0<C<\infty$), we have $\|\nabla_{\theta}\log\pi_{\theta}(y_{\text{win}}\mid x)-\nabla_{\theta}\log\pi_{\theta}(y_{\text{lose}}\mid x)\|\leq 2C$ from the triangle inequality. Taking the limit and applying Jensen’s inequality and the bounded convergence theorem, we have:
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}$ $\|\mathrm{IF}_{\mathrm{DPO}}(x,\theta,p_{\mathcal{D}})\|$
$\leq\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\bigg\|\bigg(\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}(s)\right]\bigg)^{-1}\bigg\|\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\bigg\|\mathbb{E}_{p(s_{\textrm{flip}})}\bigg[F_{\theta^{*}}(s_{\textrm{flip}})\bigg]\bigg\|$
$\leq(1/L^{\prime})\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathbb{E}_{p(s_{\textrm{flip}})}[\sigma(-g_{\theta^{*}}(s_{\textrm{flip}}))]\cdot 2C$
$=(1/L^{\prime})\cdot 1\cdot 2C=2C/L^{\prime}.$
The fact $0<2C/L^{\prime}<\infty$ according to $0<C<\infty$ and $0<L^{\prime}<\infty$ completes the proof. ∎
@@ E.3 Proof for Theorem 4
We next evaluate $\nabla_{\theta}A(\theta)=\nabla_{\theta}\mathbb{E}_{p_{\mathcal{D}}}[\sigma(g_{\theta}(s))^{\gamma}]$. Since $p_{\mathcal{D}}$ does not depend on $\theta$, from the chain rule, we can see $\nabla_{\theta}\mathbb{E}_{p_{\mathcal{D}}}[\sigma(g_{\theta}(s))^{\gamma}]=\mathbb{E}_{p_{\mathcal{D}}}[\nabla_{\theta}\sigma(g_{\theta}(s))^{\gamma}]$, where
@@
Furthermore, we show the following lemma to precisely derive the IF for Hölder-DPO.
@@
Suppose that $\theta^{*}=\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}(p_{\mathcal{D}}\|\sigma(g_{\theta}))$. Let $0<\gamma<\infty$, and let $0<\sigma(g_{\theta}(s))$. Assume that $\phi(h)$ satisfies $\phi^{\prime}(h)\neq 0$ for $h>0$ and $\phi(h)\neq c\cdot h$ for any constant $c$. Then, the first-order optimal condition of Hölder-DPO, $0=\nabla_{\theta}S_{\gamma}(p_{\mathcal{D}}\|\sigma(g_{\theta}))|_{\theta=\theta^{*}}$, holds if and only if $\mathbb{E}_{p_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}}(s)]=0$ and $\int F^{(1+\gamma)}_{\theta^{*}}(s)\mathrm{d}s=0$, where $F^{(\gamma)}_{\theta^{*}}(s)\coloneqq\sigma(g_{\theta}(s))^{\gamma}(1-\sigma(g_{\theta}(s)))\nabla_{\theta}g_{\theta}(s)$.
@@
From Proposition 2, the first-order optimal condition is given by:
$0$ $=\nabla_{\theta}S_{\gamma}(p_{\mathcal{D}}\|\sigma(g_{\theta}))\bigg|_{\theta=\theta^{*}}$
$=\underbrace{\gamma\phi^{\prime}(h{\theta})}_{C_{1}}\underbrace{\mathbb{E}{p_{\mathcal{D}}}[F^{(\gamma)}_{\theta}(s)]}_{X}+\underbrace{(1+\gamma)\left(\phi(h{\theta})-\phi^{\prime}(h_{\theta})h_{\theta}\right)}_{C_{2}}\underbrace{\bigg(\int F^{(1+\gamma)}{\theta}(s)\mathrm{d}s\bigg)}_{Y}.$
This is a linear combination $C_{1}X+C_{2}Y=0$. The “if and only if” statement holds if we can show that both coefficients $C_{1}$ and $C_{2}$ are non-zero.
First, we show $h_{\theta^{*}}>0$. By definition, $h_{\theta^{*}}=A(\theta^{*})/B(\theta^{*})$, where $A(\theta^{*})=\mathbb{E}_{p_{\mathcal{D}}}[\sigma(g_{\theta^{*}}(s))^{\gamma}]$ and $B(\theta^{*})=\int\sigma(g_{\theta^{*}}(s))^{1+\gamma}\mathrm{d}s$. From the premise $0<\sigma(g_{\theta}(s))$ and $\gamma>0$, we have $\sigma(g_{\theta^{*}}(s))^{\gamma}>0$ and $\sigma(g_{\theta^{*}}(s))^{1+\gamma}>0$. Therefore, $A(\theta^{*})>0$ and $B(\theta^{*})>0$, which implies $h_{\theta^{*}}>0$. Given $\gamma>0$, $h_{\theta^{*}}>0$, and our assumption $\phi^{\prime}(h)\neq 0$ for $h>0$, the first coefficient $C_{1}=\gamma\phi^{\prime}(h_{\theta^{*}})\neq 0$.
The second coefficient $C_{2}$ is zero if and only if $\phi(h_{\theta^{*}})-h_{\theta^{*}}\phi^{\prime}(h_{\theta^{*}})=0$. This condition holds if $\phi(h)$ is a homogeneous function of degree 1, i.e., $\phi(h)=c\cdot h$. By our assumption $\phi(h)\neq c\cdot h$, this implies $\phi(h_{\theta^{*}})-h_{\theta^{*}}\phi^{\prime}(h_{\theta^{*}})\neq 0$, and thus $C_{2}\neq 0$. Since both coefficients $C_{1}$ and $C_{2}$ are non-zero, the optimality condition $C_{1}X+C_{2}Y=0$ holds if and only if $X=\mathbb{E}_{p_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}}(s)]=0$ and $Y=\int F^{(1+\gamma)}_{\theta^{*}}(s)\mathrm{d}s=0$. This completes the proof. ∎
We remark that the conditions of $\phi(h)$, $\phi^{\prime}(h)\neq 0$ for $h>0$ and $\phi(h)\neq c\cdot h$ for any constant $c$., introduced in Lemma 1, are satisfied by the DP divergence and the PS score (which is closely related to the $\gamma$-divergence). The following lemma formally verifies this. This fact indicates that constructing $\phi$ so as to satisfy the above condition is one of keys to guarantee robustness.
@@
Let $\gamma>0$. The $\phi(h)$ functions for the DP-divergence and the PS-divergence (Remark 1) satisfy the assumptions required in Lemma 1, namely $\phi^{\prime}(h)\neq 0$ and $\phi(h)-h\phi^{\prime}(h)\neq 0$ for all $h>0$.
@@
For DP-divergence: Let $\phi(h)=\gamma-(1+\gamma)h$.
  • (i)
    The derivative is $\phi^{\prime}(h)=-(1+\gamma)$. Since $\gamma>0$, $\phi^{\prime}(h)$ is a non-zero constant, thus $\phi^{\prime}(h)\neq 0$ for all $h>0$.
  • (ii)
    We check the second coefficient term from Lemma 1:
    $\phi(h)-h\phi^{\prime}(h)$ $=\bigl(\gamma-(1+\gamma)h\bigr)-h\bigl(-(1+\gamma)\bigr)$
    $=\gamma-(1+\gamma)h+(1+\gamma)h$
    $=\gamma$
    Since $\gamma>0$, we have $\phi(h)-h\phi^{\prime}(h)\neq 0$.
For PS score: Let $\phi(h)=-h^{1+\gamma}$.
  • (i)
    The derivative is $\phi^{\prime}(h)=-(1+\gamma)h^{\gamma}$. Since $\gamma>0$ and we evaluate for $h>0$, $h^{\gamma}>0$. Thus, $\phi^{\prime}(h)$ is strictly negative, and $\phi^{\prime}(h)\neq 0$.
  • (ii)
    We check the second coefficient term:
    $\phi(h)-h\phi^{\prime}(h)$ $=\bigl(-h^{1+\gamma}\bigr)-h\bigl(-(1+\gamma)h^{\gamma}\bigr)$
    $=-h^{1+\gamma}+(1+\gamma)h^{1+\gamma}$
    $=(-1+1+\gamma)h^{1+\gamma}$
    $=\gamma h^{1+\gamma}$
    Since $\gamma>0$ and $h>0$, we have $\gamma h^{1+\gamma}>0$, which implies $\phi(h)-h\phi^{\prime}(h)\neq 0$.
In both cases, the assumptions hold. This completes the proof. ∎
@@
[…]
@@
$0$ $=\nabla_{\theta}S_{\gamma}(\tilde{p}^{(\epsilon)}_{\mathcal{D}}\|\sigma(g_{\theta}))\bigg|_{\theta=\theta^{*}(\epsilon)}$
[…]
@@
$0$ $=\gamma\bigg\{\underbrace{\frac{\partial}{\partial\epsilon}\phi^{\prime}(\tilde{h}_{\theta^{*}(\epsilon)})\mathbb{E}_{\tilde{p}^{(\epsilon)}_{\mathcal{D}}}[F^{(\gamma)}_{\theta^{*}(\epsilon)}(\tilde{s})]}_{\mathrm{(I)}}\bigg\}$
@@
[…]
@@
[…]
@@
[…]
@@
From the assumption, $(C^{(\gamma)}_{\theta^{*}}(s))^{-1}$ exists because $H_{\theta^{*}}^{(\gamma)}(s)$ is positive definite from the fact that the gradient of Hölder-DPO is $\sigma(g_{\theta^{*}}(s_{\textrm{flip}}))^{\gamma}\,\nabla_{\theta}\mathcal{L}(s_{\textrm{flip}},\pi_{\theta^{*}})$, where $\nabla_{\theta}\mathcal{L}(s_{\textrm{flip}},\pi_{\theta^{*}})$ is the gradient of DPO loss whose Hessian is assumed as positive definite. This completes the proof. ∎